OpenStudy (anonymous):

Arrange the hyperbolas in increasing order of the widths of their asymptote rectangles.

OpenStudy (anonymous):

(y-11)^2/7^2 - (x-2)^2/6^2 =1 (y-1)^2/5^2 - (x-7)^2/12^2 =1 (x-6)^2/8^2 - (y+1)^2/3^2 =1 (y-4)^2/2^2 - (x+4)^2/3^2 =1 (y+1)^2/2^2 - (x-1)^2/9^2 =1 (x+7)^2/4^2 - (y-9)^2/9^2 =1

OpenStudy (anonymous):

@aaronq @Nnesha